At a football match,there were 250 more men than women.The number of children was twice the number of women,and the number of men was twice the number of women and children combined.How many people were at the match?
Answer by ikleyn(52813)

.
At a football match,there were 250 more men than women.The number of children was twice the number of women,
and the number of men was twice the number of women and children combined.How many people were at the match?
~~~~~~~~~~~~~~~


Let x = the number of women.


Then the number of children was 2x;

the combined number of children and women was 2x + x = 3x;

and the number of men was 2*(3x) = 6x.



From the condition,


    6x - x = 250     ("there were 250 more men than women")

    5x      = 250.

     x      = 250/5 = 50


ANSWER.  50 women,  100 children and 300 men.

Solved.